Find the sum of all of the 3-digit positive integers.
Found 2 solutions by htmentor, Alan3354:
Answer by htmentor(1343)

The three digit positive integers range from 100 to 999.
This is an arithmetic sequence with a common difference of 1.
The sum of n terms of an arithmetic sequence is S_n = (n/2)(a_1 + a_n)
The last term, 999, corresponds to n = 900.
Thus S_n = (900/2)(100+999) = 494550

Answer by Alan3354(69443)

Find the sum of all of the 3-digit positive integers.
-----------
Sum from 100 to 999
---
100 to 999 --> 900 numbers
Use 450 pairs.
100 + 999 = 1099
101 + 998 = 1099
etc
450*1099 = 494550